HUBBLE SPACE TELESCOPE DAILY REPORT #3084 PERIOD COVERED: 0000Z (UTC) 03/29/02 - 0000Z (UTC) 04/01/02 Daily Status Report as of 091/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of ACS/HFC 9014 (Fine Corrector Alignment) The Advanced Camera for Surveys (HFC) was used to optimize the ACS image quality over the by adjusting the IM1 and M1 corrector mechanisms in both tip/tilt {cylinder rotation} and focus. This activity uses the same target field and exposures as proposal 9013 {ACS Coarse Corrector Alignment}, but performs fine focus and tip/tilt scans to refine the alignment resulting from that activity. The proposal completed with no reported problems. 1.2 Completed Six Sets of STIS/CCD 8963 (CCD Bias) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD read noise from multiple bias frames obtained in common science modes: amplifier D, gains 1 and 4 electrons/DN, and binning factors of 1x1, 1x2, 2x1, and 2x2. The read noise measurements before and after NCC turn-on are then compared, and standard calibration reference files for use in pipeline reduction of SMOV3B data are constructed. The proposal had no reported problems. 1.3 Completed Eleven Sets of ACS/WFC/HRC 8947 (Weekly Test) The Advanced Camera for Surveys (WFC and HRC) was used to perform basic tests to monitor, the read noise, the development of hot pixels and test for any source of noise in ACS CCD detectors. This program will be executed at least once a day for the entire lifetime of ACS. There were no problems reported. 1.4 Completed Seven Sets of WF/PC-2 8950 (SM3B Cool Down, Contamination Monitor, and Focus Check) The WF/PC-2 was used to monitor the planned cooldown to -88C, followed by UV monitors, focus checks, as well as decons and associated monitors. No problems were noted. 1.5 Completed WF/PC-2 9574 (Focus Monitor) The WF/PC-2 was used to monitor HST focus. In principle ACS/HRC should provide superior capability for monitoring the HST focus, bur since most of the HST science will shift to ACS in Cycle 11 trending observations are initiated with this program to establish the focus of WFPC2/PC1 {and hence HST} and ACS/HRC {and WFC to a lessor degree} using exposures over full orbits that are taken in parallel. No problems were reported. 1.6 Completed Five Sets of STIS/CCD 8964 (CCD Darks) The Space Telescope Imaging Spectrograph (CCD) was used to obtain dark frames twice daily, as in a corresponding calibration program. Standard calibration reference darks are constructed and hot pixel lists are updated for use in pipeline reduction of SMOV3B data. All observations completed without incident. 1.7 Completed STIS/CCD 9157 (Fundamental Properties Of L-Type Dwarfs In Binaries) The Space Telescope Imaging Spectrograph (CCD) was used to characterize the physical properties of eight L-dwarfs in four binary systems. The goal is to obtain astrometric, photometric and spectroscopic measurements of each component that will yield basic information on atmospheric and dynamical properties. The proposal completed nominally. 1.8 Completed Two Sets of FGS/1 8999 (SM3B Jitter Test) Fine Guidance Sensor 1R was used in both Position Mode {finelock} and Transfer Mode to observe selected targets in order to gain insight into the vehicle jitter characteristics after the alterations made to HST during SM3B. No problems were encountered. 1.9 Completed Two Sets of ACS/WFC/HRC 9031 (CCD Hot Pixel Annealing) The Advanced Camera for Surveys (WFC and HRC) was used to perform hot pixel annealing, which will be done once per month. The CCD TECs will be turned off and heaters will be activated to bring the WFC detector temperature to about +10C. The HRC temperature will reach about 30C. This state will be held for approximately 24 hours, after which the heaters are turned off, the TECs turned on, and the CCDs returned to normal operating condition. To assess the effectiveness of this procedure, a bias and two dark images will be taken before and after the annealing procedure for both WFC and HRC. There were no reported problems. 1.10 Completed WF/PC-2 9069 (Proper Motions in Extragalactic Optical Jets) The WF/PC-2 was used to observe the flow velocity of extragalactic jets which is a crucial missing parameter in our understanding of these objects. We would like to build on our successful HST proper motion measurements in M87, and propose similar measurements in four other optical jets. The proposal completed with no reported problems. 1.11 Completed Two Sets of FGS/1 9581 (A BEA Science Program: A High Angular Resolution Survey of Some of the Galaxy's Most Massive Stars) Fine Guidance Sensor #1 was used to make optimal use of otherwise unused telescope time while HST's pointing is restricted to the BEA cone during the first 12 days of SMOV3B. The launch date of Feb 29 placed two of the Galaxy's most active star forming regions, Trumplers 14 and 16 in the Carina Nebula, within the BEA for the entire 12 day period. FGS1r was used to conduct a high angular resolution survey of the most massive and luminous stars in these two magnificent clusters to establish the incidence of binarity at the high mass end of the HR diagram. This survey will close the detection gap between spectroscopic and speckle surveys in which duplicity can be concealed and will dramatically improve our knowledge of the IMF at the high mass end of the main sequence as well as the luminosity function {LF} of star forming regions which produce such high mass stars. Finally, this survey will identify targets for an approved SIM Key Science project to dynamically determine O-star masses. There was no anomalous activity. 1.12 Completed Four Sets of WF/PC-2 8953 (SM3B Relative Photometry Check) The WF/PC-2 was used to determine relative photometric throughput by comparing aperture photometry with pre-SMOV data. Updates to the photometric tables in the pipeline and SYNPHOT database will then be performed as needed. There were no problems with this proposal. ********************************************************** The following scheduled proposal did not execute due to the NICMOS safing at 087/1910Z: Thirty-eight Sets of NICMOS/1/2/3 8945 (Cooling System Monitoring) The following scheduled proposals did not fully execute due to the inhibition of high voltage turn-on commanded at 087/1910Z: Two Sets of STIS/MA1/MA2 8965 (MAMA Darks) STIS/CCD/MA1 8960 (Jitter Test) STIS/CCD/MA1 8962 (Contamination Monitor) 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 18 Successful: 18 Scheduled Re-acquisitions: 17 Successful: 17 2.2 FHST Updates: Scheduled: 32 Successful: 31 The roll delay update at 089/080134Z failed due to tracker #3. The subsequent full maneuver updates were successful. HSTAR 8587 was written. 2.3 Operations Notes: Operations requests were used to: Reset STIS HV event flag at 089/2231Z ACS Fine Corrector Alignment - Part 1 at 089/2229Z Delay STIS MAMA-2 turn-on at 090/0044Z The following operations notes were run over the weekend: Lim Chg JERRCNT at 088/1752Z Turn on ACS HRC/SBC corrector status limit at 088/1755Z WF/PC-2 recom at 089/0435Z to 089/1414Z PCS Gyro #2 limit change The NICMOS Cryo Cooler cool down continues nominally. SI C&DH errors were cleared at 088/1651Z and at 088/1923Z per ROP NS-5. The Zone 2B temperature (OTCZ2BT) began flagging out-of-limits intermittently at 089/040046Z. The low value was 17.8714 degC. HSTAR 8588 was written. The NSSC-1 status buffer was dumped and reset at 089/0443Z, using ROP NS-3. Per HSTAR 8586, the RGA 2 motor current (GRG1_2MC) briefly flagged out-of-limits high at 089/061634Z with a value of 128.8. The increase in current was expected due to the gyro 1 turn-off. The 486 engineering status buffer limits were adjusted at 089/0801Z as directed by ROP DF-18A. Using ROP SR-9A, the SSR-3 EDAC error counter was cleared at 089/1217Z. The SSR-1 EDAC error counter was cleared at 090/1406Z per ROP SR-1A. ESTR-2 was reconditioned at 090/1507Z per ROP RD-7A. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of Servicing Mission Orbital Verification and the gradual resumption of normal science observations and calibrations. /CAW